The amount of an invoice is $2,000, with terms 2/10, n/30. The amount to be paid within the discount period is: OA. O B. OC. OD. $1,800 $2,000. $1,400. $1,960.Lionworks Enterprises had the following inventory data: Date July 1 July 4 July 7 July 11Purchase July 14Sale Quantity Unit Cost Beginning inventory Purchase Sale 5 10 12 $50 $55 $62 8 Assuming FIFO, what is the ending inventory after the July 14 sale? (Round any intermediary calculations to the nearest cent and your final answer to the nearest dollar.) OA. $67 B. $248 OC. $210 O D. $200Which of the following is an INCORRECT statement if ending inventory is understated? 0 A. Cost of goods sold is overstated. O B. Net income understated. ° C. Income tax is understated. 0 D. Gross profit is overstated.A company uses the perpetual inventory system. At year end the general ledger indicated that this company had a balance of $50,000 in the Inventory account. Actual inventory on hand per a physical count was $50,500. What action does the company now need to take? O A. No action is needed; the difference between the ledger and actual is less O B. The company should debit the Purchases account and credit Cost of O C. The company needs to debit Inventory and credit Cost of Goods Sold for O D. The company needs to debit Cost of Goods Sold and credit Inventory, than 5%. Goods Sold $500 $500Given the following inventory activity, what is ending inventory using the perpetual average costing method? (ROund any intermedlary calculations and your tinal answer to the nearest cent.) Date Beginning Balance September 17 September 24 September 29 Quantity Unit Cost 120 60 15 20 $3.00 $4.50 Purchase Sale Purchases $6.00 OA. 165 units @ $3.50 and 20 units @ $6.00 O B. 105 units @ $3.00 and 60 units @ $4.50 and 20 units @ $6.00 O C. 120 units @ $3.00 and 45 units @ $4.50 and 20 units @ $6.00 D. 185 units @ $3.77
Answer #1

1)
Answer is D.$1960
Explanation:
The amount to be paid within the discount period = $2000 - (2000*2%) = $2000 - $40 =$1960

2)
Answer is B.$248
Explanation:
Value of ending inventory by using FIFO Methods = 4 Units *$62 = $248

3)
Answer is D. Gross profit is overstated
Explanation:
If closing inventory is understated then it overall impact would be reduced gross profit.

4)

Answer is C. The company needs to debit Inventory and Credit Cost of Goods Sold for $500.

Explanation:
Journal Entry:

Inventory account Debit $500

Cost of Goods Sold Credit $500

5)

Answer is A. 165 units @ $3.50 and 20 units @ $6.00

Explanation:
Calculation of Ending Inventory using perpetual inventory method:

Total inventory available for sale = 120 + 60 + 20 = 200 units

Total units sold = 15 units

Ending inventory = Total inventory available for sale - Total units sold = 200 units - 15 units = 185 units

Average cost per unit for before sale = 120 *3 + 60*4.50 / 120 + 60 = 360 + 270 / 180 = 630/180 = 3.50

Average cost for last purchase = $6.00 i.e. purchase price